Analytical Geometry

Two straight lines X+2Y=2 and X+2Y=6 are given ,then find the equation of the line parallel to the given lines and divided distance between the lines in the ratio 2:1 internally Ans=3X+6Y=14

aayushi verma
8 Years agoGrade

1 Answer

Harsh Patodia
8 Years ago
distance between them =4/\sqrt{5}(use formula distance between two parallel lines)
now ratio is 2:1 so distance from x+2y=2 is 8/3\sqrt{5}now again use formula between two parallel lines and you will get (if equation is x+2y=c) c=2\s\pm8/3 so c=14/3,-2/3 use 14/3 in place of c and u will get 3x+6y=14
NOTE-(distance between ax+by+c=0 and ax+by+c1=0 is c-c1/a2+b2)